# RICE: Breaking Through the Training Bottlenecks of Reinforcement Learning with Explanation

### v1

@@ 1 Introduction
Deep reinforcement learning (DRL) has shown promising performance in various applications ranging from playing simulated games (Todorov et al., 2012; Mnih et al., 2013; Oh et al., 2016) to completing real-world tasks such as navigating autonomous vehicles and performing cybersecurity attacks and defenses (Bar-Zur et al., 2023; Vyas et al., 2023; Anderson et al., 2018; Wang et al., 2023). However, training an optimal DRL agent for complex tasks, particularly in environments with sparse rewards, presents a significant challenge. Often cases, the training of a DRL agent can hit a bottleneck without making further process: its sub-optimal performance becomes evident when it makes common mistakes or falls short of achieving the final goals.
@@
Another reason behind the training bottleneck can be the poor choice of the training algorithm. Naturally, to improve performance, the developer needs to select another DRL training algorithm to re-train the DRL agent. However, for complex DRL tasks, re-training the agent from scratch is too costly. For instance, for AlphaStar (Vinyals et al., 2019) to attain grandmaster-level proficiency in StarCraft, its training period exceeds one month with TPUs. Retraining an agent of this level can incur a cost amounting to millions of dollars (Agarwal et al., 2022). Therefore, existing research has investigated the reuse of previous DRL training (as prior knowledge) to facilitate new training (Ho & Ermon, 2016; Fu et al., 2018). The most recent example is Jump-Start Reinforcement Learning (JSRL) proposed by Uchendu et al. (2023) which leverages a pre-trained policy to design a curriculum to guide the training of a self-improving exploration policy. However, their selection of exploration frontiers in the curriculum is random, which cannot guarantee that the exploration frontiers have positive returns. This motivates us to incorporate explanation methods to scrutinize the pre-trained policy and design more effective exploration frontiers.
In this work, we propose RICE, a Refining scheme for ReInforCement learning with Explanation. We first leverage a state-of-the-art explanation method to derive a step-level explanation for the pre-trained DRL policy. The explanation method identifies the most critical states (i.e., steps that contribute the most to the final reward of a trajectory), which will be used to construct the exploration frontiers. Based on the explanation results, we construct a mixed initial state distribution that combines the default initial states and the identified critical states to prevent the overfitting problem. By forcing the agent to revisit these exploration frontiers, we further incentivize the agent to explore starting from the frontiers. Through exploration, the agent is able to expand state coverage, and therefore more effectively break through the bottlenecks of reinforcement learning training. Our theoretical analysis shows that this method achieves a tighter sub-optimality bound by utilizing this mixed initial distribution (see subsection 3.4).
@@ 2.2 Leveraging Existing Policy
The utilization of existing policies to initialize RL and enhance exploration has been explored in previous literature. Some studies propose to “roll-in” with an existing policy for better exploration, as demonstrated in works (Agarwal et al., 2020; Li et al., 2023). Similar to our approach, JSRL (Uchendu et al., 2023) incorporates a guide policy for roll-in, followed by a self-improving exploration policy. However, JSRL relies on a curriculum for the gradual update of the exploration frontier. However, the curriculum may not be able to truly reflect the key reasons why the guide policy succeeds or fails. Therefore, we propose to leverage the explanation method to automatically identify crucial states, facilitating the rollout of the policy by integrating these identified states with the default initial states. In section 4, we empirically demonstrate that JSRL performs poorly in our selected games. Chang et al. (2023) propose PPO++ that reset the environment to a mixture of the default initial states and the visited states of a guide policy (i.e., a pre-trained policy). It can be viewed as a special case in our experiments, i.e., constructing a mixed initial distribution with a random explanation. However, we claim that not all visited states of a pre-trained policy are informative and our theoretical analysis and experiments both show that RICE based on our explanation method outperforms the refining method based on a random explanation.
@@ 3.1 Problem Setup and Assumption
We model the problem as a Markov Decision Process (MDP), which is defined as a tuple $<\mathcal{S}$, $\mathcal{A}$, $\mathcal{P}$, $\rho$, $\mathcal{R}$, $\gamma>$. In this tuple, $\mathcal{S}$ and $\mathcal{A}$ are the state and action set, where each $\mathit{s}_{t}$ and $\mathit{a}_{t}$ represents the state and action of the agent at time $t$. $P:\mathcal{S}\times\mathcal{A}\rightarrow\Delta(\mathcal{S})$ is the state transition function, $\mathcal{R}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the reward function. $\gamma\in(0,1)$ is the discount factor. For a policy $\pi(a|s)$: $\mathcal{S}\rightarrow\mathcal{A}$, the value function and $Q$-function is defined as $V^{\pi}(s)=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})\mid s_{0}=s\right]$ and $Q^{\pi}(s,a)=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})\mid s_{0}=s,a_{0}=a\right]$. The advantage function for the policy $\pi$ is denoted as $A^{\pi}(s,a)=Q^{\pi}(s,a)-V^{\pi}(s)$. We assume the initial state distribution is given by $\rho$: $s_{0}\sim\rho$. The goal of RL is to find an optimal policy $\pi^{*}$ that maximizes its expected total reward : $\pi^{*}=\operatorname*{arg\,max}_{\pi}\mathbb{E}_{s\sim\rho}\left[V^{\pi}(s)\right]$. Besides, we also introduce the state occupancy distribution and the state-action occupancy measure for $\pi$, denoted as $d_{\rho}^{\pi}(s)=(1-\gamma)\sum_{t=0}^{\infty}\gamma^{t}\operatorname{Pr}^{\pi}\left(s_{t}=s\mid s_{0}\sim\rho\right)$ and $d_{\rho}^{\pi}(s,a)=d_{\rho}^{\pi}(s)\pi(a|s)$.
In our setting, we have a pre-trained policy denoted as $\pi$, which may be sub-optimal. Our objective is to break through the training bottlenecks of the pre-trained policy with an explanation. Rather than re-training from scratch, we propose to utilize explanation to take full advantage of the guidance of the pre-trained policy $\pi$. Importantly, we do not assume knowledge of the original training algorithm used for policy $\pi$. And we make the following assumptions about the quality of a random policy $\pi^{r}$.
@@
Given a random policy $\pi^{r}$, we have $\mathbb{E}_{a\sim\pi^{r}}[A^{\pi}(s,a)]<=0$, $\forall s$.
@@
In other words, the pre-trained policy visits all good states in the state space. It is a standard assumption in the online policy gradient learning  (Agarwal et al., 2021; Uchendu et al., 2023). The ratio in Assumption 3.2 is also referred to the distribution mismatch coefficient.
Figure 1: Given a pre-trained DRL policy that is not fully optimal (a), we propose the RICE algorithm that reset the RL agent to specific visited states (a mixture of default initial states and identified critical states) (b), followed by an exploration step initiated from these chosen states (c).
@@ 3.3 Technique Detail
$\footnotesize a_{t}\odot a_{t}^{m}=\left\{\begin{aligned} &a_{t},&\text{if }a_{t}^{m}=0\,,\\ &a_{\text{random}}&\text{if }a_{t}^{m}=1\,,\end{aligned}\right.\vspace{-5pt}$ (1)
@@ 4 Evaluation
In this section, we start with our experiment setup and design, followed by experiment results and analysis. We provide additional evaluation details in Appendix C. For reproducibility purposes, we provide our code in the supplementary material.
@@ 4.1 Experiment Setup
Environment Selection.   We select eight representative environments to demonstrate the effectiveness of RICE across two categories: simulated games (Hopper, Walker2d, Reacher, and HalfCheetah of the MuJoCo games) and real-world applications (selfish mining, network defense, autonomous driving, and malware mutation). We additionally run the experiments in the three sparse MuJoCo games introduced by Mazoure et al. (2019). The details of these applications can be found in Appendix C.2.
@@
Baseline Refining Methods.   We compare our refining method with three baselines. The first baseline is “PPO fine-tuning” (Schulman et al., 2017), i.e., lowering the learning rate and continuing training. The second baseline is a refining method introduced by StateMask (Cheng et al., 2023), i.e., resetting to the critical state and continuing training from the critical state. The third baseline is Jump-Start Reinforcement Learning (referred to as “JSRL”) (Uchendu et al., 2023). JSRL introduces a guided policy $\pi_{g}$ to set up a curriculum to train an exploration policy $\pi_{e}$. Through initializing $\pi_{e}=\pi_{g}$, we can transform JSRL to be a refining method that can further improve the performance of the guided policy.
Evaluation Metrics.   To evaluate the fidelity of the generated explanation, we utilize an established fidelity score metric defined in StateMask (Cheng et al., 2023). The idea is to use a sliding window to step through all time steps and then choose the window with the highest average importance score (scored by the explanation method). The width of the sliding window is $l$ while the whole trajectory length is $L$. Then we randomize the action(s) at the selected critical step(s) in the selected window (i.e., masking) and measure the average reward change as $d$. Additionally, we denote the maximum possible reward change as $d_{max}$. Therefore, the fidelity score is calculated as $log(d/d_{max})-log(l/L)$. A higher fidelity score indicates higher fidelity.
For the refining methods, we measure the reward of the target agent before and after refining for all applications with dense rewards except the malware mutation application. For this application, we report the “final reward” as the probability of evading the malware detector, before and after the refining. It is worth noting that for the MuJoCo games under the sparse reward setting, we report the performance during the refining process.
@@ 4.2 Experiment Design
Experiment II.   To show the effectiveness of the refining method, we compare the agent’s performance after refining using our refining method and three aforementioned baseline methods, i.e., PPO fine-tuning (Schulman et al., 2017), StateMask’ fine-tuning from critical steps (Cheng et al., 2023), and Jump-Start Reinforcement Learning. For this experiment, all the refining methods use the same explanation generated by our explanation method if needed, to ensure a fair comparison.
@@
Experiment IV.   We test the impact of hyper-parameter choices for two primary hyper-parameters for refining: $p$ (used to control the mixed initial state distribution) and $\lambda$ (used to control the exploration bonus). For our refining method, we vary $p$ from the space $\{0,0.25,0.5,0.75,1\}$ and vary $\lambda$ from the space $\{0,0.1,0.01,0.001\}$. By examining the agent’s performance with various $\lambda$ values, we can further investigate the necessity of the exploration bonus. Additionally, we evaluate the choice of $\alpha$ for our explanation method (used to control the mask ratio for the mask network). For the explanation method, we vary $\alpha$ from the space $\{0.01,0.001,0.0001\}$.
@@ 4.3 Experiment Results
We also run our experiments of varying refining methods on selected MuJoCo games with sparse rewards. Figure 2 shows the results of our method against other baselines in SparseHopper and SparseWalker2d games. We observe that our refining method has significant advantages over other baselines with respect to final performance and refining efficiency. Through varying explanation methods, we confirm that the contribution should owe to our explanation method. We leave the refining results of the SparseHalfCheetah game and the hyper-parameter sensitivity results of all sparse MuJoCo games in Appendix C.4.
@@
First, $p$ controls the mixing ratio of critical states (identified by the explanation method) and the initial state distribution for refining. The performance is low when $p=0$ (all starting from the default initial distribution) or $p=1$ (all starting from the identified critical states). The performance has significant improvements when $p$ is between 0 and 1, i.e., using a mixed initial state distribution. Across four applications, it is more beneficial to keep $p$ between 0 and 0.5. This suggests that the probability of refining from the critical state should be lower than that from the original initial state in order to prevent overfitting.
Second, as long as $\lambda>0$ (meaning we enable the exploration), the performance has a significant improvement. The performance is less sensitive to the specific value of $\lambda$. In general, $\lambda=0.01$ has a good performance across all four applications.
Third, recall that the hyper-parameter $\alpha$ is to control the bonus of blinding the target agent. We observe that our explanation method is insensitive to the choice of $\alpha={0.01,0.001,0.0001}$.
5 Discussion and Conclusion

### v3

@@ 1 Introduction
Deep reinforcement learning (DRL) has shown promising performance in various applications ranging from playing simulated games (Todorov et al., 2012; Mnih et al., 2013; Oh et al., 2016; Cai et al., 2023) to completing real-world tasks such as navigating autonomous vehicles and performing cybersecurity attacks and defenses (Bar-Zur et al., 2023; Vyas et al., 2023; Anderson et al., 2018; Wang et al., 2023). However, training an optimal DRL agent for complex tasks, particularly in environments with sparse rewards, presents a significant challenge. Often cases, the training of a DRL agent can hit a bottleneck without making further process: its sub-optimal performance becomes evident when it makes common mistakes or falls short of achieving the final goals.
@@
Another reason behind the training bottleneck can be the poor choice of the training algorithm. Naturally, to improve performance, the developer needs to select another DRL training algorithm to re-train the DRL agent. However, for complex DRL tasks, re-training the agent from scratch is too costly. For instance, for AlphaStar (Vinyals et al., 2019) to attain grandmaster-level proficiency in StarCraft, its training period exceeds one month with TPUs. Retraining an agent of this level can incur a cost amounting to millions of dollars (Agarwal et al., 2022). Therefore, existing research has investigated the reuse of previous DRL training (as prior knowledge) to facilitate re-training (Ho & Ermon, 2016; Fu et al., 2018; Cai et al., 2022). The most recent example is Jump-Start Reinforcement Learning (JSRL) proposed by Uchendu et al. (2023) which leverages a pre-trained policy to design a curriculum to guide the training of a self-improving exploration policy. However, their selection of exploration frontiers in the curriculum is random, which cannot guarantee that the exploration frontiers have positive returns. This motivates us to incorporate explanation methods to scrutinize the pre-trained policy and design more effective exploration frontiers.
In this work, we propose RICE [^1] The source code of RICE can be found in https://github.com/chengzelei/RICE, a Refining scheme for ReInforCement learning with Explanation. We first leverage a state-of-the-art explanation method to derive a step-level explanation for the pre-trained DRL policy. The explanation method identifies the most critical states (i.e., steps that contribute the most to the final reward of a trajectory), which will be used to construct the exploration frontiers. Based on the explanation results, we construct a mixed initial state distribution that combines the default initial states and the identified critical states to prevent the overfitting problem. By forcing the agent to revisit these exploration frontiers, we further incentivize the agent to explore starting from the frontiers. Through exploration, the agent is able to expand state coverage, and therefore more effectively break through the bottlenecks of reinforcement learning training. Our theoretical analysis shows that this method achieves a tighter sub-optimality bound by utilizing this mixed initial distribution (see subsection 3.4).
@@ 2.2 Leveraging Existing Policy
The utilization of existing policies to initialize RL and enhance exploration has been explored in previous literature. Some studies propose to “roll-in” with an existing policy for better exploration, as demonstrated in works (Agarwal et al., 2020; Li et al., 2023). Similar to our approach, JSRL (Uchendu et al., 2023) incorporates a guide policy for roll-in, followed by a self-improving exploration policy. Technically, JSRL relies on a curriculum for the gradual update of the exploration frontier. However, the curriculum may not be able to truly reflect the key reasons why the guide policy succeeds or fails. Therefore, we propose to leverage the explanation method to automatically identify crucial states, facilitating the rollout of the policy by integrating these identified states with the default initial states. In section 4, we empirically demonstrate that JSRL performs poorly in our selected games. Chang et al. (2023) propose PPO++ that reset the environment to a mixture of the default initial states and the visited states of a guide policy (i.e., a pre-trained policy). It can be viewed as a special case in our framework, i.e., constructing a mixed initial distribution with a random explanation. However, we claim that not all visited states of a pre-trained policy are informative and our theoretical analysis and experiments both show that RICE based on our explanation method outperforms the refining method based on a random explanation.
@@ 3 Proposed Technique
Figure 1: Given a pre-trained DRL policy that is not fully optimal (a), we propose the RICE algorithm that resets the RL agent to specific visited states (a mixture of default initial states and identified critical states) (b), followed by an exploration step initiated from these chosen states (c).
@@ 3.1 Problem Setup and Assumption
We model the problem as a Markov Decision Process (MDP), which is defined as a tuple $\langle\mathcal{S}$, $\mathcal{A}$, $\mathcal{P}$, $\rho$, $\mathcal{R}$, $\gamma\rangle$. In this tuple, $\mathcal{S}$ and $\mathcal{A}$ are the state and action set, where each $\mathit{s}_{t}$ and $\mathit{a}_{t}$ represents the state and action of the agent at time $t$. $P:\mathcal{S}\times\mathcal{A}\rightarrow\Delta(\mathcal{S})$ is the state transition function, $\mathcal{R}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the reward function. $\gamma\in(0,1)$ is the discount factor. For a policy $\pi(a|s)$: $\mathcal{S}\rightarrow\mathcal{A}$, the value function and $Q$-function is defined as $V^{\pi}(s)=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})\mid s_{0}=s\right]$ and $Q^{\pi}(s,a)=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})\mid s_{0}=s,a_{0}=a\right]$. The advantage function for the policy $\pi$ is denoted as $A^{\pi}(s,a)=Q^{\pi}(s,a)-V^{\pi}(s)$. We assume the initial state distribution is given by $\rho$: $s_{0}\sim\rho$. The goal of RL is to find an optimal policy $\pi^{*}$ that maximizes its expected total reward : $\pi^{*}=\operatorname*{arg\,max}_{\pi}\mathbb{E}_{s\sim\rho}\left[V^{\pi}(s)\right]$. Besides, we also introduce the state occupancy distribution and the state-action occupancy measure for $\pi$, denoted as $d_{\rho}^{\pi}(s)=(1-\gamma)\sum_{t=0}^{\infty}\gamma^{t}\operatorname{Pr}^{\pi}\left(s_{t}=s\mid s_{0}\sim\rho\right)$ and $d_{\rho}^{\pi}(s,a)=d_{\rho}^{\pi}(s)\pi(a|s)$.
In our setting, we have a pre-trained policy denoted as $\pi$, which may be sub-optimal. Our objective is to break through the training bottlenecks of the pre-trained policy with an explanation. Rather than re-training from scratch, we propose to utilize explanation to take full advantage of the guidance of the pre-trained policy $\pi$. Importantly, we do not assume knowledge of the original training algorithm used for policy $\pi$. And we make the following assumptions about the quality of $\pi$.
@@
Given a random policy $\pi^{r}$, we have $\mathbb{E}_{a\sim\pi^{r}}[A^{\pi}(s,a)]\leq 0$, $\forall s$.
@@
In other words, Assumption 3.2 requires that the pre-trained policy visits all good states in the state space. Note that it is a standard assumption in the online policy gradient learning  (Agarwal et al., 2021; Uchendu et al., 2023; Li et al., 2023) and is much weaker than the single policy concentrateability coefficient assumption (Rashidinejad et al., 2021; Xie et al., 2021), which requires the pre-trained policy visits all good state-action pairs. The ratio in Assumption 3.2 is also referred to as the distribution mismatch coefficient.
@@ 3.3 Technique Detail
$a_{t}\odot a_{t}^{m}=\left\{\begin{aligned} &a_{t},&\text{if }a_{t}^{m}=0\,,\\ &a_{\text{random}}&\text{if }a_{t}^{m}=1\,,\end{aligned}\right.\vspace{-5pt}$ (1)
@@ 4 Evaluation
In this section, we start with our experiment setup and design, followed by experiment results and analysis. We provide additional evaluation details in Appendix C.
@@ 4.1 Experiment Setup
Environment Selection.   We select eight representative environments to demonstrate the effectiveness of RICE across two categories: simulated games (Hopper, Walker2d, Reacher, and HalfCheetah of the MuJoCo games) and real-world applications (selfish mining, network defense, autonomous driving, and malware mutation)  [^2] These are representative security applications that have a significant impact on the security community (Anderson et al., 2018) and they represent RL tasks with sparse rewards, which are common in security applications.. We additionally run the experiments in the three sparse MuJoCo games introduced by Mazoure et al. (2019). The details of these applications can be found in Appendix C.2.
@@
Baseline Refining Methods.   We compare our refining method with three baselines. The first baseline is “PPO fine-tuning” (Schulman et al., 2017), i.e., lowering the learning rate and continuing training with the PPO algorithm. The second baseline is a refining method introduced by StateMask (Cheng et al., 2023), i.e., resetting to the critical state and continuing training from the critical state. The third baseline is Jump-Start Reinforcement Learning (referred to as “JSRL”) (Uchendu et al., 2023). JSRL introduces a guided policy $\pi_{g}$ to set up a curriculum to train an exploration policy $\pi_{e}$. Through initializing $\pi_{e}=\pi_{g}$, we can transform JSRL to be a refining method that can further improve the performance of the guided policy.
Evaluation Metrics.   To evaluate the fidelity of the generated explanation, we utilize an established fidelity score metric defined in StateMask (Cheng et al., 2023). The idea is to use a sliding window to step through all time steps and then choose the window with the highest average importance score (scored by the explanation method). The width of the sliding window is $l$ while the whole trajectory length is $L$. Then we randomize the action(s) at the selected critical step(s) in the selected window (i.e., masking) and measure the average reward change as $d$. Additionally, we denote the maximum possible reward change as $d_{max}$. Therefore, the fidelity score is calculated as $\log(d/d_{max})-\log(l/L)$. A higher fidelity score indicates higher fidelity.
For the applications with dense rewards except the malware mutation application, we measure the reward of the target agent before and after refining. In the case of the malware mutation application, we report the “final reward” as the probability of evading the malware detector, both before and after refining. For the applications with sparse rewards, we report the performance during the refining process.
@@ 4.2 Experiment Design
Experiment II.   To show the effectiveness of the refining method, we compare the agent’s performance after refining using our method and three aforementioned baseline methods, i.e., PPO fine-tuning (Schulman et al., 2017), StateMask’s fine-tuning from critical steps (Cheng et al., 2023), and Jump-Start Reinforcement Learning (Uchendu et al., 2023). For this experiment, all the refining methods use the same explanation generated by our explanation method if needed, to ensure a fair comparison. Additionally, we conduct a qualitative study to understand how our refining method influences agent behavior and performance.
@@
Experiment IV.   To show the versatility of our method, we examine the refining performance when the pre-trained agent was trained by other algorithms such as Soft Actor-Critic (SAC) (Haarnoja et al., 2018). First, we obtain a pre-trained SAC agent and then use Generative Adversarial Imitation Learning (GAIL) (Ho & Ermon, 2016) to learn an approximated policy network. We compare the refining performance using our method against baseline methods, i.e., PPO fine-tuning (Schulman et al., 2017), StateMask’s fine-tuning from critical steps (Cheng et al., 2023), and Jump-Start Reinforcement Learning (Uchendu et al., 2023). In addition, we also include fine-tuning the pre-trained SAC agent with the SAC algorithm as a baseline.
Experiment V.   We test the impact of hyper-parameter choices for two primary hyper-parameters for refining: $p$ (used to control the mixed initial state distribution) and $\lambda$ (used to control the exploration bonus). For our refining method, we vary $p$ from $\{0,0.25,0.5,0.75,1\}$ and vary $\lambda$ from $\{0,0.1,0.01,0.001\}$. By examining the agent’s performance with various $\lambda$ values, we can further investigate the necessity of the exploration bonus. Additionally, we evaluate the choice of $\alpha$ for our explanation method (used to control the mask ratio for the mask network). Specifically, we vary $\alpha$ from $\{0.01,0.001,0.0001\}$.
@@ 4.3 Experiment Results
We also run our experiments of varying refining methods on selected MuJoCo games with sparse rewards. Figure 2 shows the results of our method against other baselines in SparseHopper and SparseHalfCheetah games. We observe that our refining method has significant advantages over other baselines with respect to final performance and refining efficiency. Through varying explanation methods, we confirm that the contribution should owe to our explanation method. We leave the refining results of the SparseWalker2d game and the hyper-parameter sensitivity results of all sparse MuJoCo games in Appendix C.4.
In addition to numerical results, we also provide a qualitative analysis of the autonomous driving case to understand how RICE influences agent behavior and performance, particularly in a critical state, in Appendix C.5. We visualize the agent’s behavior before and after refining the agent to show that RICE is able to help the agent break through the bottleneck based on the identified critical states of the failure.
@@
Refining a Pre-trained Agent of Other Algorithms.   To show that our framework is general to refine pre-trained agents that were not trained by PPO algorithms, we do experiments on refining a SAC agent in the Hopper game. Figure 3 demonstrates the advantage of our refining method against other baselines when refining a SAC agent. Additionally, we observe that fine-tuning the DRL agent with the SAC algorithm still suffers from the training bottleneck while switching to the PPO algorithm provides an opportunity to break through the bottleneck. We provide the refining curves when varying hyper-parameters $p$ and $\lambda$ in Appendix C.3.
@@
First, $p$ controls the mixing ratio of critical states (identified by the explanation method) and the initial state distribution for refining. The performance is low when $p=0$ (all starting from the default initial distribution) or $p=1$ (all starting from the identified critical states). The performance has significant improvements when $10<p<$ 1, i.e., using a mixed initial state distribution. Across all applications, we observe that setting $p$ to 0.25 or 0.5 is most beneficial. A mixed initial distribution can help eliminate the problem of overfitting.
Second, as long as $\lambda>0$ (thereby enabling exploration), there is a noticeable improvement in performance, highlighting the importance of exploration in refining the pre-trained agent. The result is less sensitive to the specific value of $\lambda$. In general, a $\lambda$ value of 0.01 yields good performance across all four applications.
Third, recall that the hyper-parameter $\alpha$ is to control the bonus of blinding the target agent when training the mask network. We vary $\alpha$ from $\{0.01,0.001,0.0001\}$ and find that our explanation method is not that sensitive to $\alpha$.
5 Discussion
Applicability.RICE is suitable for DRL applications that are trained within controllable environment (e.g., simulators), in order to generate explanations. In fact, most of today’s DRL applications rely on some form of simulator for their training. For example, for safety-critical applications such as autonomous driving, DRL agents are usually designed, trained, and tested in a simulated environment first before moving them to real-world testing. Simulation platforms broadly include Carla (Dosovitskiy et al., 2017) and MetaDrive (Li et al., 2022) which have been used to facilitate the training of DRL agents (Zhang et al., 2021; Wang et al., 2023; Peng et al., 2022). Therefore, RICE should be applicable to such DRL systems (especially during their development phase) for refining a pre-trained DRL agent.
Warm Start vs Cold Start.   As is mentioned in section 3, our method requires a “warm start” setting, i.e., the agent has good coverage of the state distribution of the optimal policy. Even if the agent has good coverage of the state distribution, it does not necessarily mean that the agent has already learned a good policy due to the potential of choosing wrong actions (Uchendu et al., 2023). Therefore, the training bottleneck can still exist under a good coverage of the state distribution. In contrast, Our method does not work well in a “cold start” setting, i.e., when the state coverage of the pre-trained policy is extremely poor. In that case, step-level explanation methods cannot give useful help and our method is actually equivalent to the RND method  [^3] We provide an example of Mountain Car game in Appendix E to illustrate this limitation..
Critical State Filtering.   Though RICE identifies critical states based on their necessity for achieving good outcomes, it does not fully consider their importance for further agent learning. For instance, a state might be deemed critical, yet the trained agent could have already converged to the optimal action for that state. In such cases, resetting the environment to this state doesn’t significantly benefit the learning process. Future work could explore additional filtering of critical states using metrics such as policy convergence or temporal difference (TD) errors, which may help concentrate efforts and accelerate refinement.
6 Conclusion